OpenStudy (anonymous):

Time spent using e-mail per session is normally distributed with mean=11 minutes and st.dev=2 minutes. a. If you select a random sample of 50 sessions, what is the probability that sample mean is between 10.8 and 11.2 minutes?

OpenStudy (anonymous):

So have you learned about "normal distributions"? Data that is distributed normally looks like a bell curve. Does that help?

OpenStudy (anonymous):

That does help but i need to find the probability I know that that is normal distribution :(

OpenStudy (anonymous):

P stands for probability P(10.8<X<11.2) = P(X<11.2) -P(X<10.8) Next, calculate the z-scores using the equation z = (X - μ) / σ For P(X<11.2), the z-score is z = (11.2-11)/2 = 0.1 Now look up the z-score using a normal distribution table, which is 0.5398. Can you take it from here? That is, can you calculate the z-score for P(X<10.8) and look it up in a normal distribution table?

OpenStudy (anonymous):

Thank you very much, now I understood the solution I can handle the continuous part. Can you also help me with another question if you don't mind?